OpenStudy (anonymous):

5x+5y=25 in slope intercept form is 5y/5=25/5-5x/5 right?

OpenStudy (anonymous):

but is the final answer y=5-1x?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

this is right .. y=5-x

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you did it :)

OpenStudy (anonymous):

im trying to figure out this problem but i think im figuring it out right have any idea how it works?

OpenStudy (anonymous):

what is x when y=0 ? what is y when x= 0 ?

OpenStudy (anonymous):

hu?

OpenStudy (anonymous):

in your question ..

OpenStudy (anonymous):

5x + 5y = 25 is correct i think so yeah because x intercept is 5 and y intercept is also 5